### v4

@@ 2.3 Contextual Semantic Integration Framework
Because SRM relies on local convolutions ($5{\times}5$ depthwise kernel) and per-position normalization, it lacks the capacity to capture long-range spatial dependencies. At the $P_{5}$ fusion level, however, defect discrimination often depends on global structural context: a damaged insulator is easier to identify when the network can reference the regular arrangement of intact neighbors, and a bird nest is distinguished from legitimate equipment by its inconsistency with the expected component layout. To provide this capability, CMAFNet inserts CSIF at the $P_{5}$ fusion stage, between C3k2 and the terminal SRM.
CSIF is placed exclusively at $P_{5}$ for two reasons. First, global attention is most effective when features encode abstract relational semantics rather than fine-grained spatial patterns. Second, the quadratic cost of self-attention ($\mathcal{O}(H^{2}W^{2})$) is tractable only at the lowest resolution ($20{\times}20$ at the nano scale); deploying CSIF at $P_{3}$ or $P_{4}$ would multiply FLOPs by $16\times$ or $4\times$, respectively, with no commensurate gain.
@@
As illustrated in Fig. 3, CSIF processes an input $\mathbf{X}\in\mathbb{R}^{B\times C_{\mathrm{in}}\times H\times W}$ through a project–refine–project pipeline. A $1{\times}1$ convolutional stem with BN and SiLU expands the channel dimension by a factor of e=2e{=}2, producing a $2C_{\mathrm{in}}$-channel tensor that is then split into two equal halves along the channel axis. One half, referred to as the attention portion, passes through a sequence of $n$ Contextual Semantic Interaction Blocks (CSIBs), while the other, referred to as the bypass portion, is forwarded without modification. After $n$ iterations, the two halves are concatenated, and a $1{\times}1$ convolution with BN and SiLU maps the result to the target dimension $C_{\mathrm{out}}$. Restricting attention to half the channels offers two advantages: it confines global interaction to a subspace of the representation, thereby mitigating the over-smoothing that full-channel attention would impose on small-object features, and it halves the computational cost of the attention operation.
@@
Each CSIB (Fig. 4) implements a standard attention–FFN cycle, with ASRM replacing conventional layer normalization in a post-normalization arrangement. The first phase computes multi-head self-attention:
@@
where $\sigma$ denotes SiLU and the hidden dimension is $2C$. A second residual connection and ASRM produce the block output: $)\mathbf{H}^{(i)}=\mathrm{ASRM}(\hat{\mathbf{U}}+\mathrm{FFN}(\hat{\mathbf{U}}$)). Post-normalization is preferred over pre-normalization because the concatenated RGB-depth features exhibit persistent distribution heterogeneity; normalizing after the residual addition corrects this distributional shift before the subsequent transformation can amplify it.
ASRM is designed to regulate the distribution heterogeneity of cross-modal fused features more flexibly than standard layer normalization. Given $\mathbf{x}\in\mathbb{R}^{B\times C\times H\times W}$, it first computes a gated interpolation between the layer-normalized representation and the original input:
@@
where $\mathrm{LN}(\cdot)$ denotes channel-wise layer normalization with statistics computed over $C$ at each spatial position, $\boldsymbol{\gamma},\boldsymbol{\gamma}_{x}\in\mathbb{R}^{C\times 1\times 1}$ are learnable per-channel scales initialized to $\boldsymbol{\gamma}{=}0.01$ and $\boldsymbol{\gamma}_{x}{=}1$, and $\odot$ denotes the Hadamard product. Under this initialization, ASRM approximates an identity mapping at the outset, thereby avoiding premature interference with the still-unstable attention weights. As training proceeds, the gate gradually shifts toward the normalized branch, effecting a smooth transition from weak to strong regulation.
@@
Because $r{=}64\ll C$, the residual formulation confines the adaptation to a low-rank subspace, ensuring that $\mathcal{F}$ acts as a correction rather than a complete feature rewrite. The multi-scale depthwise convolutions provide spatially adaptive context at three receptive-field sizes without incurring the quadratic cost of a second attention layer. This combination distinguishes ASRM from both standard gated normalization, which lacks a spatial adaptation path, and adapter-style modules, which lack the gated normalization front-end that governs the identity-to-regulation transition.
CSIF uses n=2n{=}2 CSIB blocks at $P_{5}$, striking a balance between modeling capacity and the over-smoothing that additional blocks would introduce at $20{\times}20$ resolution. Through global attention, the module enables direct interaction between RGB-derived and depth-derived features across all spatial positions, allowing the network to learn cross-modal correspondences that are beneficial for defect detection. The refined $P_{5}$ features then pass through the terminal SRM for a final round of bottleneck-based distribution normalization before entering the detection head.

### v5

@@ 2.3 Contextual Semantic Integration Framework
Because SRM relies on local convolutions ($5{\times}5$ depthwise kernel) and per-position normalization, it cannot capture long-range spatial dependencies. At the $P_{5}$ fusion level, however, defect discrimination often depends on global structural context: a damaged insulator is easier to identify when the network can reference the regular arrangement of intact neighbors, and a bird nest is distinguished from legitimate equipment by its inconsistency with the expected component layout. To provide this capability, CMAFNet inserts CSIF at the $P_{5}$ fusion stage, between C3k2 and the terminal SRM.
CSIF is placed exclusively at $P_{5}$ for two reasons. Global attention is most effective when features encode abstract relational semantics rather than fine-grained spatial patterns. In addition, the quadratic cost of self-attention ($\mathcal{O}(H^{2}W^{2})$) is tractable only at the lowest resolution ($20{\times}20$ at the nano scale); deploying CSIF at $P_{3}$ or $P_{4}$ would multiply FLOPs by $16\times$ or $4\times$, respectively, without a commensurate gain.
@@
As illustrated in Fig. 3, CSIF processes an input $\mathbf{X}\in\mathbb{R}^{B\times C_{\mathrm{in}}\times H\times W}$ through a project–refine–project pipeline. A $1{\times}1$ convolutional stem with BN and SiLU expands the channel dimension by a factor of e=2e{=}2, producing a $2C_{\mathrm{in}}$-channel tensor that is then split into two equal halves along the channel axis. One half, the attention portion, passes through a sequence of $n$ Contextual Semantic Interaction Blocks (CSIBs), while the other, the bypass portion, is forwarded without modification. After $n$ iterations, the two halves are concatenated, and a $1{\times}1$ convolution with BN and SiLU maps the result to the target dimension $C_{\mathrm{out}}$. Restricting attention to half the channels confines global interaction to a subspace of the representation, which mitigates the over-smoothing that full-channel attention would impose on small-object features and halves the computational cost of the attention operation.
@@
Each CSIB (Fig. 4) follows a standard attention–FFN cycle, with ASRM replacing conventional layer normalization in a post-normalization arrangement. The first phase computes multi-head self-attention:
@@
where $\sigma$ denotes SiLU and the hidden dimension is $2C$. A second residual connection and ASRM produce the block output: $)\mathbf{H}^{(i)}=\mathrm{ASRM}(\hat{\mathbf{U}}+\mathrm{FFN}(\hat{\mathbf{U}}$)). Post-normalization is preferred over pre-normalization because the concatenated RGB-depth features exhibit persistent distribution heterogeneity; normalizing after the residual addition corrects this distributional shift before the next transformation can amplify it.
ASRM regulates the distribution heterogeneity of cross-modal fused features more flexibly than standard layer normalization. Given $\mathbf{x}\in\mathbb{R}^{B\times C\times H\times W}$, it first computes a gated interpolation between the layer-normalized representation and the original input:
@@
where $\mathrm{LN}(\cdot)$ denotes channel-wise layer normalization with statistics computed over $C$ at each spatial position, $\boldsymbol{\gamma},\boldsymbol{\gamma}_{x}\in\mathbb{R}^{C\times 1\times 1}$ are learnable per-channel scales initialized to $\boldsymbol{\gamma}{=}0.01$ and $\boldsymbol{\gamma}_{x}{=}1$, and $\odot$ denotes the Hadamard product. Under this initialization, ASRM approximates an identity mapping at the start of training, avoiding premature interference with the still-unstable attention weights. As training proceeds, the gate gradually shifts toward the normalized branch, producing a smooth transition from weak to strong regulation.
@@
Because $r{=}64\ll C$, the residual formulation confines the adaptation to a low-rank subspace, so that $\mathcal{F}$ acts as a correction rather than a complete feature rewrite. The multi-scale depthwise convolutions supply spatially adaptive context at three receptive-field sizes without the quadratic cost of a second attention layer. This design distinguishes ASRM from standard gated normalization, which lacks a spatial adaptation path, and from adapter-style modules, which lack the gated normalization front-end that governs the identity-to-regulation transition.
CSIF uses n=2n{=}2 CSIB blocks at $P_{5}$, balancing modeling capacity against the over-smoothing that additional blocks would introduce at $20{\times}20$ resolution. Through global attention, the module allows direct interaction between RGB-derived and depth-derived features across all spatial positions, enabling the network to learn cross-modal correspondences useful for defect detection. The refined $P_{5}$ features then pass through the terminal SRM for a final round of bottleneck-based distribution normalization before entering the detection head.